# ToolACE: Winning the Points of LLM Function Calling

### v1

@@ 3 Data Verification
A critical factor influencing the function-calling capability of LLMs is the accuracy and reliability of the training data. Data that is inconsistent or inaccurate can hinder the model’s ability to interpret and execute functions [11]. Unlike general question-answering data, where verifying correctness can be challenging, function-calling data is more straightforward to validate. This is because a successful function call must strictly match the format specified in the API definition. Building on this insight, we propose an automatic dual-layer verification system (DLV) to verify our synthesized data, as shown in the right part of Figure 1, which consists of a rule verification layer, and a model verification layer, where these results are all overseen by human experts.
Figure 3: Example of the JSON format of the API definition and the function call.
Data Structure.
Each data sample contains three components: system prompt, tool list, and dialogs, all of which are stored in JSON format with the necessary pattern requirements for ToolACE. JSON format is naturally easy to parse and maintains a clear hierarchical structure. Figure 3 shows our data examples of API definition and function call. Each function call is also formatted in JSON under the field tool_usage of the assistant role, with the API name and parameters explicitly listed.
Figure 4: Examples of errors detected by the rule verification layer.
Figure 5: The pass rate of the rule verification and the model verification in DLV, respectively.
@@
The rule verification layer deploys a rule checker to ensure that the data strictly adheres to the predefined syntactic and structural requirements of the API. The quality of the data is evaluated from four key aspects: API definition clarity, function calling executability, dialog correctness, and data sample consistency, guided by a meticulously curated set of rules, as listed in Appendix A.
For instance, to verify function calling executability, we implement the following procedures: First, we confirm that the API name matches one from the given tool list. Next, we verify that all required parameters are accurately provided. Finally, we use regular expressions to ensure that the parameter formats and patterns adhere to those specified in the API documentation. These procedures allow us to validate the correctness and executability of function calls without the need for actual execution, which enhances efficiency and reduces deployment overhead. Examples of possible errors detected by the rule verification layer are displayed in Figure 5.
@@
Each aspect is evaluated by an individual expert agent, powered by an LLM. We also incorporate several other data quality verification queries to eliminate repetitive responses and meaningless tokens within the data. The pass rate of our data for the rule verification and the model verification respectively are presented in Figure 5. Additionally, we apply the self-consistency strategy mentioned in Section 2.2 with a majority voting to enhance decision accuracy.
4 Data Analysis
This section provides an in-depth analysis of our final data after verification, focusing on the distribution across multiple dimensions with respect to both diversity and complexity.
4.1 Diversity
In this section, we quantify and present the diversity of our data across three dimensions: tool diversity, format diversity, and dialog diversity. These dimensions are critical for assessing the richness of the dataset and its capacity to support robust tool learning.
4.1.1 Tool Diversity
Large API Pool.
Through our TSS module in Section 2.1, we curate a comprehensive collection of 26,507 distinct APIs, encompassing both real-world and synthesized API definitions. Our APIs span 30 first-level coarse-grained domains, 390 second-level domains and 3,398 third-level domains, significantly enriching the diversity of our tool-learning data. The left figure in Figure 6 illustrates the distribution of the 26,507 APIs across the 30 first-level domains, including science, data, and travel, among others. The right figure details the distribution of the second-level domains within the ’Entertainment’ domain. The diversity of APIs lays a robust foundation for the diversity of the overall dataset, enabling the model to generalize effectively to new APIs and tasks.
Figure 6: Left figure: Distribution statistics of first-level domains across all APIs. Right figure: Detailed distribution of second-level domains within the Entertainment domain.
Parameter Types.
We aim to maintain a balanced distribution of parameter types within the APIs. To prevent the overrepresentation of string parameters, we intentionally increase the proportion of non-string types (e.g., integer, boolean, float, array, dictionary) during the API synthesis process. The resulting distribution is depicted in Figure 8.
Tool diversity in our dataset enhances both the breadth of API coverage and the richness of parameter variety. Such diversity is crucial for the development of models that can adapt to a wide range of scenarios and inputs for real-world tasks.
4.1.2 Format Diversity
Tool Description.
As mentioned in Section 2.2.2, we store API definitions in a standard JSON format for subsequent verification. After the verification process, we convert the API definitions to multiple formats, including JSON, YAML, XML, and Markdown, to accommodate different API specification needs. Figure 17 in Appendix B illustrates examples of the diverse formats.
Function Calls.
We have created hundreds of unique function call formats through various combinations. These format requirements are explicitly declared in the system prompt, ensuring that the dialog data adheres to the specified function call formats, thereby effectively enhancing the format compliance capability of function calling. Examples of such formats are provided in Figure 9.
Format diversity strengthens the model’s ability to understand various requirements across different scenarios, thereby improving its capability to effectively follow any specific format instructions and to avoid the short cutting.
4.1.3 Dialog Diversity
Dialog Mode.
ToolACE supports a range of function call modes, including single, parallel, dependent function calls, and non-tool-use dialogs. It also covers different interaction types, such as single-turn and multi-turn dialogs. Figure 8 illustrates the distribution of these dialog modes, highlighting the extensive coverage of function call use cases in ToolACE.
Figure 7: Distribution of parameter types.
Figure 8: Distribution of data categories.
Figure 9: Function call format generalization.
4.2 Complexity
An appropriate level of complexity is essential for improving the adaptability of models to more challenging function-calling scenarios. We investigate the factors influencing complexity from two perspectives: query-level complexity and dataset-level complexity.
4.2.1 Query-level Complexity
We propose to measure the complexity of individual queries by evaluating the relevance between the user query and the description of the used tools. Generally, a higher affinity between these two elements increases the likelihood that the model will select the correct tool. Conversely, a larger discrepancy suggests that more advanced reasoning is needed to identify the appropriate function, thereby indicating a more challenging and complex query.
To quantify this relevance, we employ BGE [22] to extract embeddings and utilize cosine similarity to assess the degree of similarity. Let $e_{q}$ and $e_{t}$ be the embeddings of the query and the API description, respectively. The complexity is then defined as follows:
[…]
The distribution of the resulting complexity scores is depicted in Figure 10, with an average score of 0.439.
Figure 10: The distribution of single data complexity scores, which has been normalized to the interval [0, 1]. Scores for data that did not involve tool invocations are excluded from the visualization. The higher score represents more complexity.
4.2.2 Dataset-level Complexity
We found that basic tool capabilities, such as single function calls, require minimal data for effective model learning. In contrast, more advanced and complex capabilities, such as parallel function calls and multi-turn dialogs, usually require more training data. Therefore, we increase the proportion of complex data (i.e., parallel function calls and multi-turn dialogs) in our dataset. The distribution of these dialog modes is illustrated in Figure 8.
The distribution of dialog modes is crucial in determining the dataset’s complexity. By carefully balancing the composition of dialog modes, we can tailor the dataset’s complexity to the needs of our models, ensuring that they are sufficiently challenged to learn effectively.
@@ 5.1 Experiment Setup
To validate the effectiveness of our approach, we have conducted extensive experiments by training LLMs with the generated data. We train the open-source LLM, LLaMA3.1-8B-Instruct [2], in the supervised fine-tuning (SFT) manner. We refer to the model trained with our data as ToolACE-8B. Due to the limited resources, we adopt the parameter-efficient training strategy LoRA [9] to fine-tune the model. As for the hyper-parameters setting, we adopt one of the most common settings, which sets the rank as 16 and alpha as 32 for all modules in the model. More detailed training settings are shown in Table 2. We compare the overall performance with the state-of-the-art open-source and API-based models, including GPT-4 series  [^2] https://chatgpt.com, Gemini-series  [^3] https://gemini.google.com/ and Claude-3 series  [^4] https://www.anthropic.com/, as well as fine-tuned function calling models including Gorilla-OpenFunctions-v2 [14], xLAM-series [11], and Functionary series [12]. We then conduct in-depth ablation study to reveal the effectiveness of accuracy, diversity, and complexity.
Table 2: Hyper-parameters in experiments for training.
Learning Rate WarmUp Ratio LR Scheduler Batch Size Epochs LoRA rank LoRA alpha
$10^{-4}$ $0.1$ cosine 48 3 16 32
@@ 5.2 Overall Performance Analysis
To assess the effectiveness of the proposed ToolACE-series models regarding their functional calling capabilities, we conducted evaluations using the BFCL benchmark [23], including BFCL-v1 and BFCL-v2  [^5] The overall performances are evaluated on both and subsequent studies are evaluated on BFCL-v1 only.. This benchmark is a comprehensive and executable function call evaluation specifically designed to assess the ability of LLMs to invoke functions. The results for our ToolACE-8B model on BFCL-v1 and BFCL-v2, along with various representative models, are summarized in Table 3 and Table 4, respectively.
@@
The findings indicate that API-based models demonstrate significant advantages over open-source models, such as the Claude series and the GPT-4 series. This advantage is likely due to their larger model sizes. Open-source models fine-tuned for function calling, such as Functionary and xLAM, exhibit competitive performance, but still fall short of the leading models, especially in BFCL-v2. Our ToolACE-8B model outperforms all other API-based and open-source models in both the AST and Exec categories of BFCL-v1, and continues to exhibit substantial advantages over most models in the context of BFCL-v2, demonstrating the effectiveness of our training data for functional calling. Additionally, ToolACE excels in detecting tool relevance (or irrelevance in v2), achieving a score of 89.17%, surpassing all other models. ToolACE-8B also consistently and significantly outperforms Functionary-Small-v3.2, which is also fine-tuned from LLaMA3.1-8B-Instruct, in all categories, providing compelling evidence of its superiority.
Table 3: Performance comparison on BFCL-v1 leaderboard (updated on 08/11/2024). The top 20 models are listed for comparison.
Rank Model Overall Accuracy AST Category Exec Category Relevance
Simple Multiple Parallel
Parallel
Multiple
Simple Multiple Parallel
Parallel
Multiple
1 ToolACE-8B (FC) 91.41 89.09 95.50 92.50 90.50 98.24 94.00 90.00 85.00 89.17
2 Claude-3.5-Sonnet-0620 (Prompt) 90.53 88.55 95.00 91.50 92.50 100 96 84 80 84.17
3 Functionary-Medium-v3.1 (FC) 88.88 86.18 95.00 93.00 89.5 95.88 94.00 90.00 80.00 81.25
4 xLAM-7b-fc-r (FC) 88.76 86.36 93.50 92.00 87.50 96.47 88.00 88.00 80.00 85.00
5 GPT-4-1106-Preview (Prompt) 88.53 88.91 95.50 89.00 91.50 99.41 94.00 82.00 82.50 72.50
6 GPT-4-0613 (Prompt) 88.53 87.27 92.50 91.00 89.00 97.06 88.00 88.00 77.5 81.67
7 GPT-4-0125-Preview (Prompt) 88.12 88.55 95.00 90.00 93.00 99.41 96.00 84.00 80.00 68.33
8 Claude-3-Opus-20240229 (Prompt) 88.00 86.36 94.5 87.50 86.50 96.47 94.00 80.00 75.00 84.58
9 GPT-4o-mini-2024-07-18 (Prompt) 87.35 86.91 93.00 86.00 88.50 99.41 96.00 82.00 82.50 75.42
10 Nemotron-4-340b-instruct (Prompt) 87.18 84.36 93.50 88.50 86.50 97.65 96.00 84.00 80.00 80.42
11 GPT-4-turbo-2024-04-09 (Prompt) 87.12 87.45 96.50 91.00 90.50 99.41 96.00 82.00 77.50 64.58
12 GPT-4-1106-Preview (FC) 86.65 83.09 92.50 90.00 86.50 91.76 94.00 86.00 77.50 83.75
13 GPT-4o-2024-08-06 (FC) 85.59 82.18 90.50 93.00 84.50 80.59 90.00 86.00 77.50 87.92
14 GPT-4-0125-Preview (FC) 85.47 82.73 92.50 91.00 86.50 78.24 92.00 88.00 80.00 84.58
15 Gorilla-OpenFunctions-v2 (FC) 85.41 87.82 95.50 87.00 87.00 97.06 96.00 80.00 77.50 60.83
16 Granite-20b-FunctionCalling (FC) 85.24 82.73 90.5 85.00 81.00 87.65 90.00 86.00 80.00 88.33
17 yi-large (FC) 85.24 83.09 93.50 90.00 86.50 92.94 96.00 82.00 82.50 71.67
18 Meta-Llama-3-70B-Instruct (Prompt) 84.94 84.36 92.50 91.00 87.00 91.76 90.00 88.00 77.50 67.92
19 GPT-4-turbo-2024-04-09 (FC) 84.41 78.73 92.00 90.50 85.50 85.29 86.00 84.00 77.50 85.42
20 GPT-4o-2024-05-13 (FC) 84.12 80.55 91.00 90.00 83.00 90.59 88.00 90.00 77.50 77.08
21 Functionary-Small-v3.2 (FC) 82.82 82.73 91.00 90.00 78.50 84.12 92.00 80.00 75.00 72.92
Table 4: Performance comparison on BFCL-v2 leaderboard (updated on 08/16/2024). The top 10 models are listed for comparison.
Rank Model Overall Accuracy AST Category Exec Category Irrelevance Relevance
Simple Multiple Parallel
Parallel
Multiple
Simple Multiple Parallel
Parallel
Multiple
1 ToolACE-8B (FC) 85.77 71.03 85.78 87.13 80.17 96.86 94.00 86.00 87.50 81.44 87.80
2 GPT-4-1106-Preview (Prompt) 85.65 79.01 89.90 82.25 83.00 99.29 94.00 88.00 82.50 61.04 97.56
3 GPT-4o-mini-2024-07-18 (Prompt) 84.35 79.32 84.24 87.50 86.67 98.29 94.00 80.00 80.00 68.09 85.37
4 GPT-4-0613 (Prompt) 84.23 76.90 82.32 90.50 80.42 97.50 88.00 88.00 75.00 70.95 92.68
5 GPT-4-turbo-2024-04-09 (Prompt) 84.11 79.34 90.63 82.75 83.25 98.50 92.00 82.00 80.00 55.05 97.56
6 GPT-4-0125-Preview (Prompt) 84.09 78.61 89.30 85.62 81.42 98.50 94.00 86.00 75.00 54.93 97.56
7 Functionary-Medium-v3.1 (FC) 81.73 74.03 87.37 81.38 78.83 97.29 90.00 88.00 75.00 72.24 73.17
8 Claude-3-Sonnet-20240229 (Prompt) 80.52 71.59 79.77 79.62 82.00 93.00 92.00 88.00 80.00 48.96 90.24
9 Claude-3.5-Sonnet-20240620 (Prompt) 80.34 76.07 85.07 77.25 71.00 98.00 92.00 82.00 72.50 78.77 70.73
10 Claude-3-Opus-20240229 (Prompt) 79.75 72.25 83.07 74.50 73.42 98.79 92.00 82.00 80.00 78.03 63.41
11 GPT-4-1106-Preview (FC) 79.65 68.89 83.58 80.62 70.83 89.86 90 84 70 73.36 85.37
12 yi-large (FC) 79.53 67.84 82.68 82.5 77.33 91.86 94 82 85 68.68 63.41
13 Functionary-Small-v3.2 (FC) 79.45 68.29 81.54 77 71.42 88.29 92 86 77.5 72.02 80.49
14 GPT-4-0125-Preview (FC) 79.41 67.97 84.3 80.12 76.83 73.29 90 86 75 75.26 85.37
15 xLAM-7b-fc-r (FC) 79.36 68.43 79.3 73.88 68.5 94.21 88 88 75 80.24 78.05
16 Functionary-Small-v3.1 (FC) 78.86 70.07 83.21 83 72.92 83.93 88 84 70 68.09 85.37
17 mistral-large-2407 (FC Any) 78.82 80.59 87.42 84.25 83.25 96.86 92 86 77.5 0.29 100
18 Gorilla-OpenFunctions-v2 (FC) 78.64 72.89 79.8 78.38 66.42 94.36 92 78 72.5 64.25 87.8
19 GPT-4-turbo-2024-04-09 (FC) 78.57 63.29 83.23 83 75 81.57 88 84 75 79.47 73.17
20 Command-R-Plus (Prompt) 78.49 71.28 79.69 81.38 75.33 91.36 92 82 77.5 51.47 82.93
21 GPT-4o-2024-08-06 (FC) 78.29 69.95 80.37 84 75 77.5 90 86 72.5 81.71 65.85
@@ 5.3 Ablation Study
To further validate the effect of various mechanisms introduced in data generation and validation, we conduct a series of in-depth ablation studies from the perspective of accuracy, complexity, and diversity, respectively.
@@ 5.3.1 Ablation on Accuracy
Effects of Formalized Thinking. This part explores the impact of formalized thinking on the dialogue generation process. We randomly selected $1,000$ user queries generated by our user agent and continued the generation process under two different conditions: one with formalized thinking and the other without. We then evaluated the generated outputs using our DLV data verification module. Table 5 presents the pass rates for the two generated datasets. The results indicate that data generated with formalized thinking consistently achieves higher pass rates across both verification layers, with a significant $10\%$ absolute improvement in the final pass rate. A detailed analysis of specific errors reveals clear gaps between the two datasets in areas such as whether calling tools, tool selection, and the handling of optional parameters. These discrepancies highlight the advantages of formalized thinking in decision-making, which therefore results in the effectiveness of incorporating formalized thinking into the dialogue generation process.
Table 5: Effects of formalized thinking. “FT”: short for “Formalized Thinking”.
Method Rule-based Pass Rate Model-based Pass Rate Final Pass Rate
With FT $67.9$ $91.1$ $61.8$
W/O FT $56.1$ $88.7$ $49.8$
Effects of the verification system. As detailed in previous sections, our verification system comprises two layers: a rule checker and a model checker. To evaluate the efficacy of each layer, we train LLaMA3.1-8B-Instruct with LoRA using three distinct datasets: (1) data without any verification (denoted as w.o. dual), (2) data without model checking (denoted as w.o. model), and (3) data subjected to dual-layer verification (denoted as Final). The resulting fine-tuned models are assessed using the BFCL benchmark, with outcomes summarized in Figure 11. Comparative analysis reveals that the model trained on data without model checking surpasses that trained on unverified data in terms of both executable and overall accuracy, thereby validating the rule checker’s effectiveness. Moreover, the model trained on dually verified data significantly outperforms both ablation models in terms of AST and overall accuracy, underscoring the indispensable role of the model checker.

### v2

@@ 2.3 Dual-Layer Data Verification
A critical factor influencing the function-calling capability of LLMs is the accuracy and reliability of the training data. Data that is inconsistent or inaccurate can hinder the model’s ability to interpret and execute functions Liu et al. (2024). Unlike general question-answering data, where verifying correctness can be challenging, function-calling data is more verifiable. This is because a successful function call must strictly match the format specified in the API definition. Building on this insight, we propose an automatic dual-layer verification system (DLV) to verify our synthesized data, as shown in the right part of Figure 1, which consists of a rule verification layer, and a model verification layer, where these results are all overseen by human experts.
@@
The rule verification layer deploys a rule checker to ensure that the data strictly adheres to the predefined syntactic and structural requirements of the API, covering four key aspects: API definition clarity, function calling executability, dialog correctness, and data sample consistency, guided by a meticulously curated set of rules, as listed in Appendix B.
For instance, to verify function calling executability, we implement the following procedures: First, we confirm that the API name matches one from the given tool list. Next, we verify that all required parameters are accurately provided. Finally, we use regular expressions to ensure that the parameter formats and patterns adhere to those specified in the API documentation. These procedures allow us to validate the correctness and executability of function calls without the need for actual execution, which enhances efficiency and reduces deployment overhead.
@@
Each aspect is evaluated by an individual expert agent, powered by an LLM. We also incorporate other verification prompts to eliminate repetitive responses and meaningless tokens in the data.
@@ 3.1 Experiment Setup
To validate the effectiveness of our approach, we have conducted extensive experiments by training LLMs with the generated data. We train the open-source LLM, LLaMA3.1-8B-Instruct AI@Meta (2024), in the supervised fine-tuning (SFT) manner, for most of the experiments. We refer to the model as ToolACE-8B. We also validate our data with other backbone LLMs like Qwen-series Bai et al. (2023). Due to the limited resources, we adopt the parameter-efficient training strategy LoRA Hu et al. (2022) to fine-tune the model. As for the hyper-parameters setting, we adopt one of the most common settings, which sets the rank as 16 and alpha as 32 for all modules. We compare the overall performance with the state-of-the-art API-based and open-source models, like GPT-series  [^2] https://chatgpt.com, as well as fine-tuned function calling models including Gorilla-OpenFunctions-v2 Patil et al. (2023) and xLAM-series Liu et al. (2024). Experiments are conducted on two representative benchmarks, including BFCL Yan et al. (2024)  [^3] The overall performance is evaluated on the latest BFCL-v3 and subsequent studies are evaluated on only non-live categories since there are more testing samples in these categories, showing more robust results. and API-Bank Li et al. (2023). The two benchmarks are comprehensive and executable function call evaluations specifically designed to assess the ability of LLMs to invoke functions. We then conduct in-depth ablation study to reveal the effectiveness of accuracy, diversity, and complexity. More experimental settings including benchmark details, evaluation metrics, and training settings are shown in Appendix C.
@@ 3.2 Overall Performance Analysis
To assess the effectiveness of our ToolACE-8B model regarding its functional calling capabilities, we compare our ToolACE-8B model with various representative models. The results are summarized in Table 2 and Table 3, respectively.
@@
The findings in BFCL indicate that API-based models demonstrate significant advantages over open-source models, such as the Claude series and the GPT-4 series. Open-source models fine-tuned for function calling, such as Functionary and xLAM, exhibit competitive performance, but still fall short of the leading models. Our ToolACE-8B model outperforms most API-based and open-source models in both the AST and Exec categories of BFCL, and continues to exhibit substantial advantages over all the open-source models in the context of API-Bank, demonstrating the effectiveness of our training data for functional calling. This is mainly attributed to our accurate, diverse, and complex synthesized data, which enhances the zero-shot function calling capability of the LLM. Additionally, ToolACE excels in mitigating hallucination, achieving impressive relevance and irrelevance scores of 85.37% and 83.81%, respectively. These results highlight its ability in maintaining an excellent balance between the two categories, unlike other models that either suffer from significant imbalance or underperform in both categories. ToolACE-8B also consistently and significantly outperforms xLAM-7b-fc-r, which is also fine-tuned for function calling with similar size, in all categories, providing compelling evidence of its superiority. Furthermore, our ToolACE-8B model shows consistent advantageous performance on API-Bank compared with all open-source models, demonstrating comparable performance with GPT-4-series models.
Table 2: Accuracy performance comparison on BFCL-v3 leaderboard (updated on 09/20/2024). The top 20 models are listed for comparison. FC denotes the model is tailored for functional calling. (A) and (E) present AST and executable category, respectively. Rel and Irrel are abbreviations for relevance and irrelevance.
Rank Overall Model Single turn Multi turn Hallucination
Non-live (A) Non-live (E) Live (A) Multi turn Rel Irrel
1 59.49 GPT-4-turbo-2024-04-09 (FC) 82.65 83.80 73.39 21.62 70.73 79.79
2 59.29 GPT-4o-2024-08-06 (FC) 85.52 82.96 71.79 21.25 63.41 82.91
3 59.22 ToolACE-8B (FC) 89.27 90.07 73.21 14.37 85.37 83.81
4 59.13 xLAM-8x22b-r (FC) 89.75 89.32 72.81 15.62 97.56 75.23
5 58.45 GPT-4o-mini-2024-07-18 (FC) 82.83 81.80 67.53 25.75 82.93 71.83
6 57.94 xLAM-8x7b-r (FC) 88.44 85.89 71.97 15.75 92.68 72.35
7 57.21 GPT-4o-mini-2024-07-18 (Prompt) 86.54 87.95 72.77 11.62 80.49 79.20
8 55.82 mistral-large-2407 (FC) 84.12 83.09 67.17 20.50 78.05 48.93
9 55.67 GPT-4-turbo-2024-04-09 (Prompt) 91.31 88.12 67.97 10.62 82.93 61.82
10 54.83 Claude-3.5-Sonnet-20240620 (FC) 70.35 66.34 71.39 23.50 63.41 75.91
11 53.66 GPT-4o-2024-08-06 (Prompt) 80.90 77.89 73.88 6.12 53.66 89.56
12 53.43 GPT-4o1-mini-2024-09-12 (Prompt) 75.48 76.86 71.17 11.00 46.34 88.07
13 53.01 Gemini-1.5-Flash-Preview-0514 (FC) 77.10 71.23 71.17 13.12 60.98 76.15
14 52.53 Gemini-1.5-Pro-Preview-0514 (FC) 75.54 77.46 69.26 10.87 60.98 80.56
15 51.93 GPT-3.5-Turbo-0125 (FC) 84.52 81.66 59.00 19.12 97.56 35.83
16 51.78 FireFunction-v2 (FC) 85.71 84.23 61.71 11.62 87.80 52.94
17 51.78 Open-Mistral-Nemo-2407 (FC) 80.98 81.46 61.44 14.25 65.85 59.14
18 51.45 xLAM-7b-fc-r (FC) 86.83 85.02 68.81 0.00 80.49 79.76
19 51.01 Gorilla-OpenFunctions-v2 (FC) 87.29 84.96 68.59 0.00 85.37 73.13
20 49.63 Claude-3-Opus-20240229 (FC) 58.40 63.16 70.50 15.62 73.17 76.40
21 49.55 Meta-Llama-3-70B-Instruct (Prompt) 87.21 87.41 63.39 1.12 92.68 50.63
Table 3: Accuracy performance comparison on API-Bank evaluation system. Bold values represent the highest performance for API-based and open-source models, respectively.
Model Call Retrieval+Call
gpt-3.5-turbo-0125 70.43 52.59
gpt-4-0613 75.94 48.89
gpt-4-turbo-2024-04-09 72.43 39.26
gpt-4o-mini-2024-07-18 74.69 45.93
API-based gpt-4o-2024-05-13 76.19 42.96
Alpaca-7B 24.06 5.19
ChatGLM-6B 23.62 13.33
Lynx-7B 49.87 30.37
xLAM-7b-fc-r 32.83 21.48
Open-source LLaMA-3.1-8B-Instruct 71.18 37.04
ToolACE-8B 75.94 47.41
@@ 3.3.1 Ablation on Accuracy
Effects of the verification system. As detailed in previous sections, our verification system comprises two layers: a rule checker and a model checker. To evaluate the efficacy of each layer, we train LLaMA3.1-8B-Instruct with LoRA using three distinct datasets: (1) data without any verification (denoted as w.o. dual), (2) data without model checking (denoted as w.o. model), and (3) data subjected to dual-layer verification (denoted as Final). It is important to note that datasets with more verification layers contain smaller amounts of data, as some data is filtered out during the verification process. The resulting fine-tuned models are assessed using the BFCL benchmark, with outcomes summarized in Figure 3. Comparative analysis reveals that the model trained on data without model checking surpasses that trained on unverified data in terms of both executable and overall accuracy, thereby validating the rule checker’s effectiveness. Moreover, the model trained on dually verified data significantly outperforms both ablation models in terms of AST and overall accuracy, underscoring the indispensable role of the model checker.